25 ° C. An aqueous solution of iodine containing 0.0516 g liter at temperature 4: 412 g liter containing CCL, is in equilibrium with the solution. 25 C If the solubility of iodine in water at temperature is 0.34 g liter-1, multiply its solubility in cc.


1
Expert's answer
2021-04-02T05:12:12-0400

At 25º C an aqueous solution of iodine containing 0.0516 g/litre–1 is in equilibrium with a carbon tetrachloride (CCl4) solution containing 4.412 g/litre–1. The solubility of iodine in water at 25 ºC is 0.34 g/litre–1. Find the solubility of iodine in carbon tetrachloride.


Solution:

First we find the value of K:

K = "\dfrac{C_{CCl_4}}{C_{H_2O}}" = "\frac{4.412}{0.0516}" = 85.5

To find the solubility of iodine in CCl4 we should use a distribution law:

"\dfrac{Solubility \ of \ iodine \ in \ CCl_4}{Solubility \ of \ iodine \ in \ water}=85.5"


Solubility of iodine in CCl4 = 85.5 "\times" 0.34 = 29.07 g / L.


Answer: 29.07 g / L.